OpenStudy (anonymous):

Approximately 2% of the nation’s children (about 1.7 million) have a parent who is in jail or prison. Let X be the number of children that have an incarcerated parent from a random sample of 100 children. a) Verify that this is a binomial setting and write it in notation. b) Describe what P(X=0) means in context. c) Find P(X=0) and P(X=1). d) What is the probability that two or more of the children have a parent in jail or prison?

OpenStudy (anonymous):

@amistre64 @experimentX

OpenStudy (anonymous):

@Agent_Sniffles

OpenStudy (anonymous):

im betting $ 50 on @campbell_st :)

OpenStudy (anonymous):

anyone?

OpenStudy (campbell_st):

well it appears to be binomial probability so using the formula \[P(X) = ^nC_{r} q^{n - r}p^r\] (a) so the the probability that a parent is in jail is 0.02 = p probability that a parent is not in jail is 0.98 = q the sum of binomial probabilities must be 1. 0.02 + 0.98 = 1 so the notation it would be and the value of n = 100.... since 100 children in the survey \[^{100}C_{r}(0.98)^{100 - r}(0.02)^r\] (b) P(X = 0) will be the probability that out of 100 children none will have a parent in jail. so r = 0 (c) to find P(X = 0) and P(X = 1) substitute r = 0 and evaluate then r = 1 and evaluate (d) for this question use the complement 1 - pron no parent in jail - prob 1 parent in jail P(x >=2) = 1 - P(X = 0) - P(X = 1) its a while since I've done binomial probability but I think this is close.

OpenStudy (anonymous):

@satellite73 plz help?

OpenStudy (anonymous):

@marsss

OpenStudy (anonymous):

i am not familiar. so sorry :)

OpenStudy (anonymous):

its ok do u know this:

OpenStudy (anonymous):

A new antibiotic is effective for 85% of infections. The antibiotic is given to 40 patients. a) Verify that this is a binomial setting and write it in notation. b) Describe what P(X=35) means in context. c) What is the probability that the antibiotic will work in at least 35 of the 40 patients? d) What is the probability that the antibiotic will work in less than half of the patients?

OpenStudy (anonymous):

@Luis_Rivera do u know this one^